OpenStudy (anonymous):

Each computation can be completed more efficiently by using an algebraic strategy rather than direct calculation. Try to identify that strategy and use it to complete each each problem. Show work. a) (42)(38) b) 543^2-321^2/543+321 c) 75^2+2(75)(25)+25^2

OpenStudy (whpalmer4):

Here's an example that might apply to one of them: \[(x-a)(x+a) = x^2-a^2\]

OpenStudy (anonymous):

Does (x-a)(x+a)= x^2- a^2 applies to the first one?

OpenStudy (whpalmer4):

Well, observe that 42 and 38 are the same distance away from 40...